OpenStudy (tabdo1226):

Mr. Evans paid $455 for a 7-day car rental. At that rate, how much would he pay for a 10-day rental? A. $65 B. $465 C. $650 D. $462

OpenStudy (tabdo1226):

@texaschic101

OpenStudy (automaticloveletter):

455/7=65 so 65 dollars each day 65*3=195 (to make 10 days) 455+195=650

OpenStudy (tabdo1226):

thanks

OpenStudy (automaticloveletter):

no problem :)

OpenStudy (tabdo1226):

Which rate is the best deal for a package of peanuts? A. $1.40 for 50 g B. $2.05 for 75 g C. $2.25 for 80 g D. $1.02 for 35 g

OpenStudy (tabdo1226):

@automaticloveletter